### v1

# Risk Aware Safe Control with Cooperative Sensing for Dynamic Obstacle Avoidance

Abstract
This paper presents the design, development, and on-vehicle implementation and validation of a safety-critical controller for autonomous driving under sensing and communication uncertainty. Cooperative sensing, fused via a Wasserstein barycenter (WB), is used to optimize the distribution of the dynamic obstacle locations. The Conditional Value-at-Risk (CVaR) is introduced to form a risk-aware control-barrier-function (CBF) framework with the optimized distribution samplings. The proposed WB–CVaR–CBF safety filter improves control inputs that minimize tail risk while certifying forward invariance of the safe set. A model predictive controller (MPC) performs path tracking, and the safety filter modulates the nominal control inputs to enforce risk-aware constraints. We detail the software architecture and integration with vehicle actuation and cooperative sensing. The approach is evaluated on a full-scale autonomous vehicle (AV) in scenarios with measurement noise, communication perturbations, and input disturbances, and is compared against a baseline MPC–CBF design. Results demonstrate improved safety margins and robustness, highlighting the practicality of deploying the risk-aware safety filter on an actual AV.
keywords:
Safe Navigation, Intelligent autonomous vehicles, Control Barrier Function
thanks: This research was supported by the CARMEN+ University Transportation Center, sponsored by the U.S. Department of Transportation under Grant No. 69A3552348327. The views presented are those of the authors and do not necessarily represent the official views of the U.S. Department of Transportation.
1 Introduction
Safety is a crucial aspect of developing autonomous vehicles (AVs). AVs utilize a variety of sensors to enhance their perception systems, enabling them to reconstruct the environment, identify vulnerable road users (VRUs), and improve decision-making and control policy. The work shown in Zhang et al. (2023) examines data fusion algorithms from multiple sensors, primarily LiDAR, cameras, and radar, for object detection to enhance autonomous driving capabilities. Also, Vehicle-to-Everything (V2X) communication is a key feature in level 4 and 5 AVs and is included as one of the sensors in Wang et al. (2020). V2X technology can extend the perception range and provide valuable insights into complex traffic environments. In Xiang et al. (2023), the authors discuss the recent research regarding multi-sensor fusion and cooperative perception in autonomous driving. The paper discusses AVs that utilize multi-sources and V2X information in cooperative perception to reconstruct the environment.
A controlled system can leverage quadratic programming or nonlinear programming to guarantee safety by defining a safe set. Then, the Control Barrier Functions (CBFs) are added as constraints to ensure that the system remains in the safe set through the principle of set-invariance (Ames et al., 2019). CBFs have been widely adopted in safety-critical systems Black et al. (2022). The discrete-time CBF (DCBF) integrated with MPC was proposed by Zeng et al. (2021), where the authors evaluate the MPC-DCBF in a racing car overtaking scenario.
However, uncertainties in vehicle and obstacle localization pose significant challenges to real-world autonomous driving. Communication delays in V2X-based sensor fusion can introduce positional biases, thereby increasing the risk of unsafe maneuvers and potential accidents. Additionally, the behavior of VRUs is characterized by an indeterminate dynamic model and inherent uncertainty. In Nair et al. (2022), the authors examine dynamics using three different distributions and apply these as constraints within three policies: robust MPC, stochastic MPC, and distributionally robust (DR) MPC. Similarly, the authors in Long et al. (2024) address uncertainty in CBF formulations by incorporating probability-based uncertainty to design a safety-critical control. They introduce the conditional value at risk (CVaR) into the DR-CBF framework to establish a safety strategy for robots. Additionally, the authors in Zhang et al. (2024) utilize CVaR within the CBF framework for vehicle overtaking scenarios, thereby providing a courteous control strategy and developing a risk map to aid in decision-making. Renganathan (2025) introduces a CVaR-CBF controller to assure safety for AVs and Chang et al. (2025) proposes a risk-budgeted monitor framework to switch the controller between MPC-CBF and CVaR controller.
The development of CBF and MPC frameworks has transitioned from deterministic to robust and stochastic frameworks that account for real-world uncertainties. The initial works in CBF assumed that the controller works with a perfect state estimation Ames et al. (2019); Zeng et al. (2021); Black et al. (2022). Following the initial work, the literature saw the introduction of measurement-robust CBFs (MR-CBFs) Cosner et al. (2021) and high-order MR-CBFs (HO-MR-CBFs) Oruganti et al. (2023) for addressing safe performance under measurement errors. The recent literature addresses CBFs in a stochastic (chance-constrained) Li et al. (2023) and distributionally robust setting (DR-CBF) Long et al. (2024). The DR-CBF addresses the problem of uncertainties in the environment that are introduced due to sensor and perception delays. Under practical implementations, these delays could cause significant safety issues in dynamic environments where AVs operate.
It is also significant that most of the literature fails to address the safety performance of CBFs under measurement noise and input disturbances. The power of sensor fusion for the reduction of inherent uncertainty is not fully harnessed. As shown in Fig. 1(c), uncertainty in the relative state between the vehicle and the dynamic obstacle can induce unsafe maneuvers, potentially leading to a collision. Thus, to address this gap, we propose a risk-aware safe controller for autonomous driving applications that builds on top of DR-CBF. The proposed method uses the Wasserstein Barycenter (WB) to utilize cooperative sensing data to provide an optimal probability distribution of the obstacle position. Within the optimization framework, the risk-aware obstacle avoidance is performed by optimizing conditional value-at-risk (CVaR) from the obstacle position distribution. The proposed WB-CVaR-CBF is integrated into the MPC framework, which ensures both trajectory following and safety guarantees in an uncertain environment. Specifically, we make the following contributions:
  • We mathematically formulate the framework to use the Wasserstein Barycenter for cooperative sensing and to derive optimal probabilistic distribution.
  • We then demonstrate how to utilize this optimal probabilistic measure to minimize the CVaR and formulate the Wasserstein Barycenter-Conditional Value-at-Risk-Control Barrier Function (WB-CVaR-CBF), and provide its feasibility criterion.
  • Finally, we show the on-vehicle implementation on a full-scale AV for dynamic obstacle avoidance, and benchmark the proposed WB–CVaR–CBF safety layer against standard CBF baselines. We release the ROS2 based software and experiment artifacts to allow reproducibility at the repository. [^1] Github repository: https://github.com/OSU-CAR-MSL/Risk-Aware-Safety-Critical-Control-for-AVs-via-Wasserstein-Barycenter-WB-CVaR-CBF-
Figure 1: (A) The proposed framework with multiple sensors and wasserstein barycenter with CVaR-CBF (B) The experimental vehicle and the sensors mounted position (C) Overview of the problem statement depicting uncertainty in vehicle position and obstacle position.
2 Preliminaries and Problem formulation
A nonlinear continuous time and control affine system is given by
$\dot{\mathbf{x}}=f(\mathbf{x})+g(\mathbf{x})\mathbf{u},$ (1)
where $\mathbf{x}\in D\subset\mathbb{R}^{n}$ and $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represent the system state and control input. $\mathcal{U}$ is the set of admissible control inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},\quad$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are locally Lipschitz continuous.
2.1 Safe Sets and Control Barrier Functions
For the given control affine system in (1), a safe set $\mathcal{C}$ is defined as the superlevel set of a continuously differentiable function $h(\mathbf{x})$, which can be written as:
$\mathcal{C}=\{\mathbf{x}\in\mathcal{D}\subset\mathbb{R}^{n}:h(\mathbf{x})\geq 0\}.$ (2)
The set $\mathcal{C}$ is rendered forward invariant by a controller $\mathbf{u}\in\mathcal{U}$ if for every $\mathbf{x_{0}}\in\mathcal{C}$, $\mathbf{x_{t}}\in\mathcal{C}$ for and all $t>t_{0}$.
The system is safe if we can implement the control input that would render the set $\mathcal{C}$ forward invariant. From Ames et al. (2019); Oruganti et al. (2024), for a given set $\mathcal{C}$, $h(\mathbf{x})$ is a control barrier function (CBF) with $\frac{\partial h}{\partial(\mathbf{x})}(\mathbf{x})\neq 0$ if there exists an extended class $\mathcal{K}_{\infty}$ function $\alpha$ such that for the control system (1):
$\sup_{\mathbf{u}\in\mathcal{U}}\left[L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})\mathbf{u}+\alpha(h(\mathbf{x}))\right]\geq 0,\quad\forall\mathbf{x}\in\mathcal{C}.$ (3)
$L_{f}h(\mathbf{x})$ is the is the Lie derivative of $h$ along $f$, and $L_{g}h(\mathbf{x})$ is the is the Lie derivative of $h$ along $g$.
The formulation in (4) is referred to as the Control Barrier Function-based Quadratic Program (CBF-QP) that ensures safety.
$\mathbf{u}=$ $\underset{\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{m}}{\text{min}}\;\frac{1}{2}\|\mathbf{u}-\mathbf{u}_{\text{nom}}\|^{2}$ (4)
s.t. $L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})\mathbf{u}+\alpha\left(h(\mathbf{x})\right)\geq 0,$
where $\mathbf{u}_{\text{nom}}$ is the nominal control input derived from the nominal controller, which is MPC in this work.
2.2 Problem Formulation
The uncertainties addressed in this paper can be divided into two primary sources: (i) the uncertainty in AV’s position. This could be due to input disturbances or the localization measurement uncertainty from the AV’s GPS and (ii) the measurement uncertainties from cooperative sensing through multi-sources. This section explores how uncertainty is modeled. We consider obstacle position measurements obtained from cooperative sensing through three types of sensors: LiDAR, camera, and V2X. All measurement positions are represented in the East-North-Up (ENU) coordinate system, which serves as the inertial reference frame.
2.2.1 AV Localization Uncertainty:
We denote the true vehicle’s $x,y$ position at time step $k$ as $\mathbf{Z}_{v,k}=(x_{v,k},y_{v,k})$, and model the measured position as a Gaussian random variable:
$\hat{\mathbf{Z}}_{v,k}\sim\mathcal{N}(\mathbf{Z}_{v,k}+\mu^{(\text{GPS})},\sigma^{({\text{GPS}})}),$ (5)
where $\hat{\mathbf{Z}}_{v,k}\in\mathbb{R}^{2}$ is the measured position, $\mu^{(\text{GPS})}$ is the mean deviation and $\sigma^{({\text{GPS}})}$ is a diagonal covariance matrix representing the position uncertainty.
2.2.2 Cooperative sensing Measurement Uncertainty:
The dynamic model of the obstacle is assumed to be as follows:
$\mathbf{o_{s,k+1}}=f_{o}(\mathbf{o_{s,k}})+\omega_{s,k},$ (6)
where $s\in\{\text{LiDAR},\text{Camera},\text{V2X}\}$, $\mathbf{o_{s,k}}\subset\mathbb{R}^{n}$ denotes the states of the obstacle at time k, $f_{o}$ is an unknown function, and $\omega_{s,k}$ is disturbance regarding to the sensors. The true obstacle position is denoted as $\mathbf{Z}_{o,k}=(x_{o,k},y_{o,k})\in\mathbb{R}^{2}$, and the obstacle position measurement from the sensor is denoted as $\hat{\mathbf{Z}}_{o,k}^{(s)}=(x_{o,k},y_{o,k})\in\mathbb{R}^{2}$ at time step $k$.
In this setup, all three sensors perceive the same dynamic obstacle, but each may include different noise characteristics due to various factors. For instance, LiDAR sensors can produce distorted point clouds due to multipath effects or highly reflective surfaces, which may lead to the inaccurate location of obstacles. Camera-based systems are prone to motion blur during fast vehicle movement, mainly when the frame rate is low (e.g., below 30 Hz), degrading image quality and affecting the performance of vision-based detection or tracking algorithms. V2X communication relies on wireless transmission, which is susceptible to latency and packet loss, especially in dynamic environments. All of these reasons can lead to outdated position information, introducing inconsistency in cooperative sensing.
In order to model the noise distribution, we assume that the measurement uncertainty in each sensor follows a Gaussian distribution on Hilbert spaces, with sensor-specific covariance $\sigma^{(s)}$, and mean bias denoted as $\mu^{(\text{s})}$. Due to V2X latency and network-induced delays, we consider the possibility that the measurements is biased, especially for fast-moving or dynamic obstacles. The resulting Gaussian distribution models for the sensors are as follows:
$\hat{\mathbf{Z}}_{o,k}^{(\text{LiDAR})}$ $\sim\mathcal{N}(\mathbf{Z}_{o,k}+\mu^{(\text{LiDAR})},\sigma^{(\text{LiDAR})}),$ (7a)
$\hat{\mathbf{Z}}_{o,k}^{(\text{Camera})}$ $\sim\mathcal{N}(\mathbf{Z}_{o,k}+\mu^{(\text{Camera})},\sigma^{(\text{Camera})}),$ (7b)
$\hat{\mathbf{Z}}_{o,k}^{(\text{V2X})}$ $\sim\mathcal{N}(\mathbf{Z}_{o,k}+\mu^{(\text{V2X})},\sigma^{(\text{V2X})}).$ (7c)
Let $L^{2}(\mathcal{D})$ denote the space of square-integrable functions mapping from $\mathcal{D}$ to $\mathbb{R}^{n}$. The covariance function $\sigma^{(s)}$ is associated with an integral operator $\Sigma^{(s)}:L^{2}(\mathcal{D})\to L^{2}(\mathcal{D})$Mallasto and Feragen (2017). The covariance operators for three sensors can be written as $\Sigma^{(\text{LiDAR})}$, $\Sigma^{(\text{Camera})}$, $\Sigma^{(\text{V2X})}$.
3 Cooperative sensing and Wasserstein Barycenter
This section discusses on how Wasserstein Barycenter is formulated for cooperative sensing. For defining the Wasserstein metric and Barycenters in the Wasserstein space, consider two vectors $\xi_{1}$ and $\xi_{2}$, supported on a set $\Xi\subseteq\mathbb{R}^{n}$ associated with probability measures $\mu^{(1)}$ and $\mu^{(2)}$, respectively. Let $\mathcal{P}_{p}(\Xi)\subseteq\mathcal{P}(\Xi)$ be the Borel probability measures over $\Xi$ with finite $p$-order moments. With $p\geq 1$, the $p$-Wasserstein distance between the probability measures $\mu^{(1)}$ and $\mu^{(2)}\in\mathcal{P}_{p}(\Xi)$ is given by
$W_{p}(\mu^{(1)},\mu^{(2)}):=\left(\inf_{\pi\in\Pi(\mu^{(1)},\mu^{(2)})}\int_{\Xi\times\Xi}\|\xi_{1}-\xi_{2}\|^{p}\mathrm{d}\pi(\xi_{1},\xi_{2})\right)^{\frac{1}{p}},$ (8)
where $\|.\|$ is the Euclidean norm on $\mathbb{R}^{n}$, $\Pi(\mu^{(1)},\mu^{(2)})$ is the set of joint distribution on $\Xi\times\Xi$ with marginals $\mu^{(1)}$ and $\mu^{(2)}$. $p=2$ is the Lévy–Fréchet metric, commonly used for comparing distributions. Intuitively, consider $\mu^{(1)}$ and $\mu^{(2)}$ to measure the same quantity using two different sensors. However, with respect to our problem formulation, we have three sensors (camera, LiDAR, and V2X) measuring the same obstacle, following three distributions - $(\mu^{(\text{LiDAR})},\mu^{(\text{Camera})},$ and $\mu^{(\text{V2X})})$. Thus, the mean of the three distributions can be determined through the Wasserstein Barycenter Agueh and Carlier (2011).
From Agueh and Carlier (2011), the $\lambda$-weighted empirical 2-Wasserstein Barycenter for finite set of probability measures $\{\mu^{(1)},\mu^{(2)},...,\mu^{(N)}\}$ with second moments is defined by
$\mu^{(\text{WB})}=\inf_{\mu}\sum_{s=1}^{N}\lambda_{s}W_{2}^{2}(\mu,\mu^{(s)}),$ (9)
where $\lambda_{s}$ are positive weights such that $\sum_{s=1}^{N}\lambda_{s}=1$. In our current formulation, the 2-Wasserstein Barycenter is considered for its existence and uniqueness properties.
3.1 Example on our three distribution
In this paper, we consider the measurement distribution of sensors to be Gaussian. Thus, we have
Lemma 1
Mallasto and Feragen (2017) Let $\{\hat{\mathbf{Z}}_{o,k}^{(s)}\}_{s=1}^{N}$ be a Gaussian measurement with $\hat{\mathbf{Z}}_{o,k}^{(s)}\sim\mathcal{N}(\mathbf{Z}_{o,k}^{(s)}+\mu^{(\text{s})},\Sigma^{(s)})$, then there exists a unique barycenter $\hat{\mathbf{Z}}_{o,k}^{(\text{WB})}\sim\mathcal{N}({\mathbf{Z}_{o,k}+\mu^{(\text{WB})}},{\Sigma^{(\text{WB})}})$ with barycentric coordinates $\{\lambda_{s}\}_{s=1}^{N}$.
If $\hat{\mathbf{Z}}_{o,k}^{(\text{WB})}$ is non-degenerate, then $\mathbf{Z}_{o,k}+\mu^{(\text{WB})}$} and $\Sigma^{(\text{WB})}$} satisfy
${\mathbf{Z}_{o,k}+\mu^{(\text{WB})}}=\sum_{s=1}^{N}\lambda_{s}(\mathbf{Z}_{o,k}^{(s)}+\mu^{(\text{s})}),$ (10a)
${\Sigma^{(\text{WB})}}=\sum_{s=1}^{N}\lambda_{s}\left({\Sigma^{(\text{WB})}}\Sigma^{(s)}{\Sigma^{(\text{WB})}}\right)^{1/2}.$ (10b)
Each weight of the sensors is defined as a positive scalar: $\lambda_{1}^{(\text{LiDAR})}$, $\lambda_{2}^{(\text{Camera})}$, $\lambda_{3}^{(\text{V2X})}$, with $\sum_{s=1}^{3}\lambda_{s}=1$. Substituting (7) and the weights into (10) gives:
$\mathbf{Z}_{o,k}+\mu^{(\text{WB})}$ $=\lambda_{1}^{(\text{LiDAR})}\left(\mathbf{Z}_{o,k}^{(\text{LiDAR})}+\mu^{(\text{LiDAR})}\right)$
$\quad+\lambda_{2}^{(\text{Camera})}\left(\mathbf{Z}_{o,k}^{(\text{Camera})}+\mu^{(\text{Camera})}\right)$
$\quad+\lambda_{3}^{(\text{V2X})}\left(\mathbf{Z}_{o,k}^{(\text{V2X})}+\mu^{(\text{V2X})}\right)$ (11a)
$\Sigma^{(\text{WB})}$ $=\lambda_{1}^{(\text{LiDAR})}\left(\Sigma^{(\text{WB})}\Sigma^{(\text{LiDAR})}\Sigma^{(\text{WB})}\right)^{1/2}$
$\quad+\lambda_{2}^{(\text{Camera})}\left(\Sigma^{(\text{WB})}\Sigma^{(\text{Camera})}\Sigma^{(\text{WB})}\right)^{1/2}$
$\quad+\lambda_{3}^{(\text{V2X})}\left(\Sigma^{(\text{WB})}\Sigma^{(\text{V2X})}\Sigma^{(\text{WB})}\right)^{1/2}$ (11b)
The obstacle position based on Wasserstein Barycenter from all sensors can be written as:
$\hat{\mathbf{Z}}_{o,k}^{(\text{WB})}\sim\mathcal{N}(\mathbf{Z}_{o,k}+\mu^{(\text{WB})},\sigma^{(\text{WB})}),$ (12)
where $\sigma^{\text{WB}}$ is the covariance of Wasserstein Barycenter.
3.2 Wasserstein Barycenter-based Control Barrier Function
In this subsection, we present the integration of the Wasserstein Barycenter into the construction of Control Barrier Functions (CBFs). The CBF at time step $k$ is defined as:
$h({\mathbf{Z}}_{v,k},{\mathbf{Z}}_{o,k})$ $=\left\|{\mathbf{Z}}_{v,k}-{\mathbf{Z}}_{o,k}\right\|_{2}-D,$ (13a)
$D$ $=R_{v}+R_{o}+d_{\text{s}},$ (13b)
where $R_{v}$ and $R_{o}$ denote the radius of the AV and the obstacle, and $d_{\text{s}}$ is the safety distance between the AV and the obstacle. (13a) considers the true positions of both the vehicle and the obstacle. However, in practical scenarios, only noisy state measurements and estimates are available. It should also be noted that the CBF should be formulated from two Gaussian measurements of the vehicle position and the Wasserstein Barycenter. Hence, (13a) is reformulated using $I$ samples from the measurement distribution in (5) and with $J$ samples from the Wasserstein Barycenter distribution in (12). Thus, the CBF formulation from sampling the two distributions is given by:
$h_{I\times J}(\hat{\mathbf{Z}}_{v,k},\hat{\mathbf{Z}}_{o,k}^{\text{(WB)}})=\left\|\hat{\mathbf{Z}}_{(v,k),i}-\hat{\mathbf{Z}}_{(o,k),j}^{\text{(WB)}}\right\|_{2}+D,$ (14a)
$\quad i=1,\ldots,I,\text{and}\quad j=1,\ldots,J.$
For notational simplicity, we denote the CBF in (14a) as $h(\cdot)$. The modified CBF optimization constraint based on sampling the Wasserstein Barycenter obstacle position and the vehicle position is given by
$\text{CBC}\left(\hat{\mathbf{Z}}_{v,k},\mathbf{u},\hat{\mathbf{Z}}_{o,k}^{\text{(WB)}}\right)=L_{f}h(\cdot)+L_{g}h(\cdot)\mathbf{u}+\alpha\left(h(\cdot)\right)\geq 0.$ (14b)
Note that the CBC in (14b) follows a distribution $\mathbb{P}$ and the next section discusses on how to compute a safe control action on this distribution.
4 Risk-Aware Safety-Critical Control Formulation
Due to the stochasticity in obstacle and vehicle position, the CBF in optimization formulation is considered a chance-constrained optimization problem that accommodates uncertainties. Thus, the chance-constraint is given by
$\mathbb{P}\left(\text{CBC}\left(\hat{\mathbf{Z}}_{v,k},\mathbf{u},\hat{\mathbf{Z}}_{o,k}^{\text{(WB)}}\right)\geq 0\right)\geq 1-\epsilon,$ (15)
where $\epsilon\in(0,1)$ is a user defined risk tolerance. For example, $\epsilon=0.05$ denotes $5\%$ risk tolerance. Note that this does not mean that the system is under 5% risk, but implies that the $\epsilon$-percentile risk is optimized. For notional simplicity, from here on, the CBC in (14b) is denoted as $CBC(.)$. The constraint in (15) makes the optimization formulation non-convex and computationally intractable Nemirovski and Shapiro (2007).
However, Rockafellar et al. (2000) shows a convex CVaR approximation of the chance constraint. We formulate the risk-based definitions as follows: Definition 3: The Value at Risk (VaR) for the CBC with distribution $\mathbb{P}$ at a risk-level $\epsilon$ for $\epsilon\in(0,1)$ is defined as
$\text{VaR}_{\epsilon}^{\mathbb{P}}\left(CBC(.)\right):=\inf_{\gamma\in\mathbb{R}}\{\gamma\>|\>\left(\mathbb{P}\left(CBC(.)\right)\leq\gamma\right)\geq\epsilon\}.$
The VaR is the inverse of the Cumulative distribution function (CDF). Clearly, $\gamma\in\mathbb{R}$ implies that it is a value in loss and does not provide much information about the distribution tail and still could make the optimization intractable. Alternatively, one could use CVaR, which is defined by Definition 4.
Definition 4: The Conditional Value at Risk (CVaR) for the CBC with distribution $\mathbb{P}$ at risk-level $\epsilon$ for $\epsilon\in(0,1)$ is defined as the
$\text{CVaR}_{\epsilon}^{\mathbb{P}}\left(CBC(.)\right)\triangleq$
$\mathbb{E}_{\mathbb{P}}\left[CBC(.)\;\middle|\right.\left.CBC(.)\leq\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))\right].$
The CVaR can be seen as the stochastic safety violation of the autonomous vehicle with VRUs. In the following part, it is shown that the CVaR can be written as a tractable convex function.
Lemma 2
Rockafellar et al. (2000) The approximated function, sampled from the distribution of CBC, given by
$F_{\epsilon}\big(\mathrm{CBC}(\cdot),\gamma\big)=\gamma-\frac{1}{N\epsilon}\sum_{i=1}^{N}\big[\gamma-\mathrm{CBC}(\cdot)\big]^{+},$ (16)
is convex.
{pf}
The proof follows from Rockafellar et al. (2000). Clearly,
$\frac{1}{N}\sum_{i=1}^{N}\big[\gamma-\mathrm{CBC}(\cdot)\big]^{+}=\mathbb{E}_{P}\!\left[\,\gamma-\mathrm{CBC}(\cdot)\,\right]^{+},$
and
$\big[\gamma-\mathrm{CBC}(\cdot)\big]^{+}=\max\{0,\,\gamma-\mathrm{CBC}(\cdot)\}.$
Since $\max\{0,\,x\}$ is convex and $\gamma$ is constant, $\max\{0,\,\gamma-\mathrm{CBC}(\cdot)\}$ is convex whenever $\mathrm{CBC}(\cdot)$ is convex. Expectations preserve convexity, hence (16) is convex.
Theorem 1
The CVaR of the CBC is determined by
$\min_{\gamma\in\mathbb{R}}F_{\epsilon}\big(CBC(.),\gamma\big).$ (17)
{pf}
The proof is inspired from Rockafellar et al. (2000). From the definition of VaR and Lemma 2, it is evident that the values of $\gamma$ that give the minimum of $F_{\epsilon}\big(CBC(.),\gamma\big)$, is $\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))$. Thus,
$\min_{\gamma\in\mathbb{R}}$ $\;F_{\epsilon}\big(CBC(.),\gamma\big)=F_{\epsilon}\big(CBC(.),\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))\big)$
$=$ $\;\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))+$
$\frac{1}{N\epsilon}\sum_{i=1}^{N}\big[CBC(.)-\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))\big]^{+}.$
From the definition of CVaR, this can be approximated as
$\approx\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))+\text{CVaR}_{\epsilon}^{\mathbb{P}}(CBC(.))$
$-\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))=\text{CVaR}_{\epsilon}^{\mathbb{P}}(CBC(.)).$
Corollary 1 Minimizing the safety violations is equivalent to minimizing $F_{\epsilon}\big(CBC(.),\gamma\big)$ over all $(\mathbf{u},\gamma)\in\mathcal{U}\times\mathbb{R}$
$\min_{\mathbf{u}\in\mathcal{U}}\text{CVaR}_{\epsilon}^{\mathbb{P}}\left(CBC(.)\right)=\min_{(\mathbf{u},\gamma)\in\mathcal{U}\times\mathbb{R}}F_{\epsilon}\big(CBC(.),\gamma\big).$ (18)
  • $F_{\epsilon}\big(CBC(.),\gamma\big)$ can be minimized jointly over the variables $(\mathbf{u},\gamma)$ through convex optimization
  • The pair $(\mathbf{u}^{*},\gamma^{*})$ achieves joint minimum iff $\mathbf{u^{*}}$ minimizes CVaR and $\gamma^{*}$ is the VaR associated with $\mathbf{u^{*}}$.
Corollary 1 intuitively shows that it is much simpler to work on the convex formulation (16) with respect to $(\mathbf{u},\gamma)$ than the direct optimization of VaR, which could be intractable. The definitions used for VaR, CVaR, and the above shown proofs differ from the literature as here we directly identify the value at risk based on the $\epsilon$-percentile of the distribution. This implies that the following constraints are equivalent.
$\mathbb{P}\{CBC(.)\leq\gamma\}\geq\epsilon\leftrightarrow\mathbb{P}\{CBC(.)\geq\gamma\}\geq 1-\epsilon$
$\leftrightarrow\text{VaR}_{\epsilon}^{\mathbb{P}}(CBC(.))\leq\gamma.$
In short, we identify the tail of the distribution that contributes to the worst-case risk and formulate the optimization problem to minimize it.
4.1 WB-CVaR-CBF-OP Formulation
This subsection shows the optimization problem (OP) formulation: Wasserstein Barycenter - Conditional Value at Risk - Control Barrier Function Optimization Problem (WB-CVaR-CBF-OP) and discusses the feasibility of the OP. Inspired from the QP formulation (4), we have
$\mathbf{u}_{\text{safe}}=$ $\underset{\mathbf{u}\in\mathcal{U},\,\gamma\in\mathbb{R}}{\text{min}}\;\frac{1}{2}\left\|\mathbf{u}-\mathbf{u}_{\text{nom}}\right\|^{2},$ (19)
$\text{s.t.}\;\gamma$- $\frac{1}{N\epsilon}\sum_{i=1}^{N}\left[\gamma-\text{CBC}\left(\hat{\mathbf{Z}}_{v,k},\mathbf{u},\hat{\mathbf{Z}}_{o,k}^{\text{(WB)}}\right)\right]^{+}\geq 0.$
(WB-CVaR-CBF-OP)
In the remainder of this section, we discuss the specific assumptions under which this OP formulation works.
Assumption 1: The camera, LiDAR, and V2X identify and track the same object. This assumption is quite valid from the perspective of the controller, that the sensor fusion, sensing, and tracking algorithms provide this information Huang et al. (2023).
The feasibility of our WB-CVaR-CBF OP formulation is implied through Assumption 1, Lemma 1, Lemma 2, and Theorem 1. Lemma 1 shows that a unique barycenter always exists. So, with Assumption 1, the control barrier constraint (14b) always exists. The optimization problem in (19) is feasible through convex optimization, as shown in Corollary 1. So, the solution is feasible as long as $CBC(.)$ is convex.
The optimization problem can become infeasible when the AV is very close to the obstacle. This can be the case for late detections, suddenly appearing obstacles, and AVs or obstacles at high speeds, which are rare and extreme scenarios and are not considered in this work.
5 Experimental and Numerical Example: Autonomous Vehicle Control
5.1 Experimental Setup
The proposed framework is assessed using both experimental trials and numerical simulations. Experimentally, we demonstrate on an autonomous vehicle (AV) that the controller can be deployed in real time and successfully avoid moving obstacles. The test vehicle is shown in Fig. 1, and the experimental environment follows the protocol in Renganathan et al. (2025). To evaluate robustness under varying noise levels, we performed Monte Carlo simulations with 100 runs per scenario. In both the experimental and simulation settings, the AV and vulnerable road users (VRUs) follow predefined trajectories that intersect such that the VRU enters the AV’s path within a 10 m radius. The AV is expected to detect the VRU and execute a safe collision-avoidance maneuver.
5.1.1 Scenarios and Parameters
Evaluations primarily focus on the controller’s performance under varying levels of uncertainty. Specifically, we validate the controller in the following three scenarios:
  • Scenario 1: The AV receives low-uncertainty GPS data, and all three sensors operate under low-noise conditions.
  • Scenario 2: The AV receives high-uncertainty GPS data, while the remaining sensors still operate with low noise.
  • Scenario 3: Both GPS and V2X measurements are affected by high levels of uncertainty and noise.
Both the optimization problems are solved using IPOPT solver in CasADi framework under the same set of parameters as shown in table 1 to ensure the comparability of the results.
Table 1: Parameters
Vehicle parameter and Constraints Value
Safety Distance $3m$
Vehicle Radius $1.8m$
Obstacle (VRU) Radius $1m$
Obstacle (VRU) Speed $4.5m/s$
$\lambda_{\text{LiDAR}},\lambda_{\text{Camera}},\lambda_{\text{V2X}}$ $0.4,0.4,0.2$
5.1.2 Comparative Analysis
For comparison, all trajectory-following controllers use MPC as the nominal controller and differ only in the safety filter:
  • CBF: We enforce the CBF constraint in (4) along the entire trajectory, using the mean of the three sensor measurements as the filter input. The CBF is computed using this mean position at each time step $k$, given by:
    $h({\hat{\mathbf{Z}}}_{v,k},{\hat{\mathbf{Z}}}_{o,k}^{\text{(mean)}})=\left\|\hat{\mathbf{Z}}_{v,k}-\hat{\mathbf{Z}}_{o,k}^{(\text{mean})}\right\|_{2}-D,$ (20)
    where $\hat{\mathbf{Z}}_{o,k}^{(\text{mean})}$ is the mean position from all the sensors measurement.
5.1.3 Metrics
To evaluate performance over all $N_{t}$ test cases, we compute the following metrics. For all metrics except the success rate, averages are taken only over the successful runs, $N_{s}$.
  • Success Rate (SR): A run is deemed successful if the minimum vehicle–pedestrian separation along the trajectory is always greater than the required clearance $d_{\min}=2.8\,\mathrm{m}$ (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$). The success rate is $\mathrm{SR}=N_{s}/N_{t}$.
  • Minimum Distance to Pedestrian (MDP): For each successful run, let $d_{m,i}$ denote the minimum distance to the nearest pedestrian during avoidance. We report the mean over successful runs,
    $\mathrm{MDP}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}d_{m,i}.$
    An ideal clearance of at least $5.8\,\mathrm{m}$ (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$ + $3.0\,\mathrm{m}$ safety buffer) satisfies the desired safety margin. Values below $5.8\,\mathrm{m}$ indicate behavior that prioritizes reference-path tracking over conservative separation.
5.2 Implementation Results
The controller stack is implemented in Python and deployed as a set of ROS 2 nodes (Fig. 2). sensing nodes (Camera, LiDAR, V2X, and GNSS) publish raw and preprocessed measurements on ROS 2 topics. A fusion node computes the Wasserstein barycenter of these sensor distributions to obtain a unified estimate of obstacle states, which is then re-published for downstream modules. The nominal trajectory-following command $\mathbf{u}_{\text{nom}}$ is generated by an MPC node. A safety-filter node subscribes to the fused obstacle state, and at each control cycle solves a QP with CVaR–CBF constraints to minimally modify $\mathbf{u}_{\text{nom}}$, producing the safe command $\mathbf{u}_{\text{safe}}=\{\text{acceleration},\text{steering}\}$. The resulting control input is passed to a CAN bus node and transmitted to the vehicle actuators via the CAN bus. All inter-module communication is handled through ROS 2 topics and services.
Figure 2: On-vehicle software architecture and data flow.
The trajectory under Scenario 2 using the proposed controller is shown in Fig. 3. A numerical simulation with the same sensor noise setting is provided in Fig. 4 scenario 2. These results demonstrate that the proposed control framework is deployable on a full-scale autonomous vehicle and can successfully avoid a moving obstacle. The remaining discrepancy in tracking performance between experiment and simulation primarily stems from the bicycle model vehicle dynamics used in the simulator.
Figure 3: On-vehicle trajectory with MPC-WB-CVaR-CBF under scenario 2.
Figure 4: Trajectories under MPC–WB–CVaR–CBF (proposed) and MPC–CBF–QP (baseline). (a) Scenario 1: both safe; (b) Scenario 2: proposed safe, baseline unsafe; (c) Scenario 3: proposed safe, baseline unsafe.
5.3 Numerical Testing Results
Having validated on-vehicle implementation, we assess robustness to sensor noise through a Monte Carlo study. The next section details 100 runs numerical experiments for each scenario.
5.3.1 Scenario 1:
This scenario has low GPS uncertainty and low obstacle detection noise, and this is close to an ideal scenario. The GPS operates with $\sigma^{(\text{GPS})}=0.1$, LiDAR with $\sigma^{(\text{LiDAR})}=0.1$, camera with $\sigma^{(\text{Camera})}=0.2$, and $\sigma^{(\text{V2X})}=1.0$. The V2X is perceived to have higher noise due to transmission delays. The results for this case are summarized in table-2. In this setting, both MPC-CBF and MPC-WB-CVaR-CBF reach 100% safety in all trials. The trajectory figure comparison has been shown in Fig. 4 (a).
Table 2: Comparison with pedestrian interactions over 100 runs in scenario 1; $\mu^{(\text{GPS})}=0,\ \sigma^{(\text{GPS})}=0.1$.
Method $\mu^{(\mathrm{LiDAR})}=0,\ \sigma^{(\mathrm{LiDAR})}=0.1;$ $\mu^{(\mathrm{Camera})}=0,\ \sigma^{(\mathrm{Camera})}=0.2;\\ \mu^{(\mathrm{V2X})}=0,\ \sigma^{(\mathrm{V2X})}=1.0$.
SR (%) MDP (m)
CBF 100% 6.15
WB-CVaR-CBF 100% 7.03
5.3.2 Scenario 2:
In this scenario, GPS uncertainty is attributed to a highly disturbed environment, such as GPS attenuation. The GPS noise level is assumed to be $\sigma^{(\text{GPS})}=0.5$, corresponding to a 0.5-meter error. Meanwhile, the obstacle positions detected by the sensors remain at a low noise level, as discussed in the previous scenario. The results are presented in table 3. Under this condition, the MPC-CBF controller encounters a 33% unsafe behavior, while the proposed MPC-WB-CVaR-CBF achieves significantly better performance with only 3% unsafe trials. Fig. 4(b) show differences in the resulting trajectories, with the proposed controller exhibiting superior safety relative to the baseline.
Table 3: Comparison with pedestrian interactions over 100 runs in scenario 2; $\mu^{(\text{GPS})}=0,\ \sigma^{(\text{GPS})}=0.5$.
Method $\mu^{(\mathrm{LiDAR})}=0,\ \sigma^{(\mathrm{LiDAR})}=0.1;$ $\mu^{(\mathrm{Camera})}=0,\ \sigma^{(\mathrm{Camera})}=0.2;\\ \mu^{(\mathrm{V2X})}=0,\ \sigma^{(\mathrm{V2X})}=1.0$.
SR (%) MDP (m)
CBF 67% 6.43
WB-CVaR-CBF 97% 6.93
Table 4: Comparison with pedestrian interactions over 100 runs in scenario 3; $\mu^{(\text{GPS})}=0,\ \sigma^{(\text{GPS})}=0.5$.
Method $\mu^{(\mathrm{LiDAR})}=0,\ \sigma^{(\mathrm{LiDAR})}=0.1;$ $\mu^{(\mathrm{Camera})}=0,\ \sigma^{(\mathrm{Camera})}=0.2;\\ \mu^{(\mathrm{V2X})}=-1,\ \sigma^{(\mathrm{V2X})}=1.0$.
SR (%) MDP (m)
CBF 92% 6.535
WB-CVaR-CBF 100% 7.168
5.3.3 Scenario 3:
This scenario considers extreme cases in V2X, where the V2X signals are significantly affected by factors such as transmission delays, network congestion, or other disturbances. The GPS uncertainty levels remain the same as in Scenario 2. Thus, we have an input disturbance (through GPS) and measurement noise through V2X. The V2X measurement distribution is characterized by a bias and variance of $\mu^{(\text{V2X})}=-1$ and $\sigma^{(\text{V2X})}=0.5$, respectively. The obstacle detections from LiDAR and camera sensors maintain low noise, as in the previous cases. The results for this scenario are shown in table 4 and Fig. 4 (a). Under these conditions, the proposed MPC-WB-CVaR-CBF controller achieves a 100% safe rate across all trials, while the MPC-CBF controller still results in an 8% unsafe rate.
6 Conclusion
We demonstrated a on-vehicle implementation of an MPC–WB–CVaR–CBF safety controller that maintains safe navigation under localization and cooperative-sensing uncertainty. The paper provides an implementation-ready framework to encode sensor trust and communication quality via distributional fusion, and exposes CVaR parameter as a practical safety performance hyperparameter for deployment.
Three uncertain scenarios are evaluated using Monte Carlo approach— (1) low variance Gaussian noise in GPS and cooperative sensing, (2) high GPS uncertainty with low variance cooperative sensing, and (3) biased, noisy V2X atop GPS uncertainty. These simulation studies consistently showed improved safety margins and robustness over a baseline MPC–CBF. The on-vehicle experiment (VRU crossing) corroborated these trends, demonstrating that the deployed WB–CVaR–CBF safety layer can be practically integrated with MPC to maintain the autonomous driving maneuver.